### v2

# Refining the grading of irreducible Lie colour algebra representations

Abstract.
We apply the loop module construction of Mazorchuk–Zhao in the context of Lie colour algebras. We construct a bijection between the equivalence classes of all finite-dimensional graded irreducible Lie colour algebra representations from the irreducible representations for Lie superalgebras. This bijection is obtained by applying the loop module construction iteratively to simple groups in the Jordan–Hölder decomposition of the grading group. Restricting to simple groups in this way greatly simplifies the construction. Despite the bijection between Lie colour algebra representations and Lie superalgebra representations, Lie colour algebras maintain a non-trivial representation theory distinct from that of Lie superalgebras. We demonstrate the applicability of the loop module construction to Lie colour algebras in two examples: a Hilbert space for a quantum mechanical model and representations of a colour version of $\mathfrak{sl}_{2}$.
Key words and phrases:
Color Lie (super)algebras, Graded Lie (super)algebras, irreducible representations
2020 Mathematics Subject Classification:
17B75, 17B70, 17B10
1. Introduction
Lie colour algebras are a generalisation of Lie superalgebras to grading by any abelian group $\Gamma$ (rather than just $\mathbb{Z}_{2}$) and were introduced by Rittenberg and Wyler [1, 2] in 1978 (however, see also [3]). Lie colour algebras initially found applications in (for example): de Sitter spaces [4, 5, 6], quasispin [7], strings [8] and extensions of Poincaré algebras [9, 10]. There has been a recent increase in activity surrounding applications of $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$-graded Lie colour algebras. This recent activity has mainly been focussed on parastatistics [11, 12, 13, 14, 15, 16, 17, 18, 19, 20] and $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$-graded quantum mechanics [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32].
Many Lie colour algebra applications make use of irreducible representations to construct various models. A classification of such representations would therefore be useful. Indeed, in [33], the authors conclude their paper with a desire for a classification of all irreducible representations for the $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$-supersymmetry algebra with $\mathcal{N}\geq 2$. Such a classification motivates the results of this paper.
Various classifications of the algebras themselves under different assumptions have already been performed [34, 35, 36, 37, 38]. Moreover, the representation theories of the colour versions of the following algebras have been examined: $\mathfrak{gl}$, $\mathfrak{sl}$, $\mathfrak{osp}$ [39, 40], the Lie algebra for the group of plane motions [41], the Heisenberg Lie algebra [42], super Schrödinger algebras [43] and a supersymmetry algebra [33].
One of the most powerful techniques in the study of Lie colour algebras is that of discolouration, which provides a bijection between Lie colour algebras and $\Gamma$-graded Lie superalgebras. This bijection preserves much of the important structure, such as subalgebras, ideals and subrepresentations. Discolouration, in the specific case of $\mathbb{Z}_{2}^{n}$-gradings, was introduced concurrently with the algebras themselves [1] and was soon generalised [44] to grading by any finitely generated abelian group.
Discolouration was reformulated in [45], using so-called Klein operators lying outside the universal enveloping algebra. It has since been shown that these Klein operators can be expressed in an algebraic extension of the universal enveloping algebras for a colour version of $\mathfrak{gl}$ [46] or for the non-coloured supersymmetric $\mathfrak{gl}(m|n)$ [47] and that Klein operators appear in the $q\to-1$ limit of $U_{q}(\mathfrak{sl}_{3})$ [48]. This version of discolouration has already been applied to the study of the representation theory. For example, in [47] it is shown how the covariant representations of $\mathfrak{gl}(m|n)$ can be lifted to the colour version of $\mathfrak{gl}$.
Despite widespread awareness of the results of [44], discolouration has rarely been utilised in the recent literature surrounding applications of $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$-graded Lie colour algebras. Perhaps a partial explanation lies with the following nuance: the discolouration of a Lie colour algebra representation is a $\Gamma$-graded representation, whose representation theory can be quite different from the more familiar $\mathbb{Z}_{2}$-graded representation theory of Lie superalgebras. For example, a $\Gamma$-graded irreducible representation may become reducible when considering only the $\mathbb{Z}_{2}$-grade.
Concurrently (yet largely independently) gradings on (non-colour) Lie algebras have been systematically studied, beginning with [49]. The classification of fine gradings on finite-dimensional simple Lie algebras over an algebraically closed field with characteristic zero has been completed—see the monograph [50] and references therein, together with [51, 52, 53]. Gradings over fields with nonzero characteristic (see e.g. [50]) and over non-algebraically closed fields (see e.g. [54]) are also under study.
Given that some Lie colour algebras discolour to give graded Lie algebras (i.e. graded Lie superalgebras with trivial odd sector), the classification results of [50, 51, 52, 53] for gradings on Lie algebras are also useful for classifying Lie colour algebras. Similarly, a classification of gradings on Lie superalgebras would aid in classifying Lie colour algebras (see [55] and references therein for work in the direction of classifying Lie superalgebra gradings). Application of the grading classification results to the problem of classifying Lie colour algebras has been studied in [36].
Although the research into graded algebras has mainly focussed on classifying the fine gradings, the representation theory of graded algebras has also received attention. Various works have shown how one can construct the graded irreducible modules from ungraded modules via different methods: induced modules [56], thin coverings [57] or loop modules [58].
In this paper, we will focus on the loop module construction, since loop modules are a natural structure to consider in the study of Lie colour algebras. Indeed, loop constructions have already appeared in the Lie colour algebra literature, often under a different name or left unnamed. For example, the ‘enhancement’ of a Lie colour algebra defined in [59] is a twisted version of the loop algebra studied in [60]. Additionally, we will show (Section 3.1) that the $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$-graded quantum mechanical Hilbert space realisation in [23] can be constructed as a loop module.
The initial motivation for introducing loop modules was to study graded irreducible representations: every $\Gamma$-graded irreducible representation is the loop module of some ungraded irreducible representation [58]. Moreover, the loop module can be used to construct an equivalence between two categories (see [61, Theorem 4.16] for details) and a bijection between equivalence classes of $\Gamma$-graded irreducible and $\Gamma/H$-graded irreducible modules [61, Remark 7.2]. Although these results were initially proved in the context graded (non-colour) Lie algebras, the results still hold for Lie colour algebras. We make extensive use of these results throughout this paper.
The goal of this paper is twofold: (1) to apply the loop module construction to Lie colour algebras and demonstrate its utility in this context; and (2) to present a more constructive loop module procedure to simplify computations.
For goal (1), we have already mentioned that the loop module already appears in the Lie colour algebra literature through enhancements [59] and a $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$-graded quantum mechanical Hilbert space [23]. In addition, we apply the loop module construction (from goal (2)) to a colour version of $\mathfrak{sl}_{2}$ to derive all of its irreducible modules. We compare these irreducible colour $\mathfrak{sl}_{2}$ modules with those that have already appeared in the literature [62, 63].
The key idea for goal 2 is to take iterated loop modules, only increasing the size of the grading group by a finite simple group $H$ at each step. This simplifies the construction due to the following result: for a finite-dimensional $\Gamma/H$-graded irreducible module $V$, either $V$ or its loop module (but not both) is a $\Gamma$-graded irreducible module. If the desired grading group is finite (which is often the case in applications of Lie colour algebras) we can use the Jordan–Hölder decomposition of the grading group to build up to the full grading group.
This iterated loop module process provides a method for deriving every finite-dimensional $\Gamma$-graded irreducible representation from the corresponding $\mathbb{Z}_{2}$-graded representations. In particular, we obtain bijections $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$ that map between equivalence classes (with equivalence relations stronger than isomorphism) of irreducible representations, as per the following diagram.
[…]
These bijections are provided in Theorem 4.8 (the existence of similar bijections were noted in [61, Remark 7.2]). The $\mathbb{Z}_{2}$-graded irreducible superalgebra representations have been well-studied, and these bijections will allow us to carry across these results to Lie colour algebras. It is important to note that these bijections are more complicated than discolouration and allow for interesting differences in the representation theories.
This paper is organised as follows. In Section 2, we review some known results for Lie colour algebras, and results analogous to those for Lie (super)algebras. We then summarise the loop module construction in Section 3. As an example, Section 3.1 uses the loop module to construct the Hilbert space of the $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$-graded supersymmetric system of [23]. In Section 4, we show how an iterated loop module construction can be used to derive the finite-dimensional $\Gamma$-graded irreducible modules from the $\Gamma/H$-graded modules. We give an example of deriving the finite-dimensional irreducible modules of a colour version of $\mathfrak{sl}_{2}$ (Section 5) and compare with [62, 63]. We end with some concluding remarks in Section 6.
2. Lie colour algebras and their representations
A Lie colour algebra generalises a Lie superalgebra to grading by some (additive) abelian group $\Gamma$ (instead of just $\mathbb{Z}_{2}$). For the formal definition of a Lie colour algebra, we refer the reader to [44]. As an illustrative example: if $A=\bigoplus_{\alpha\in\Gamma}A_{\alpha}$ is an associative algebra such that $A_{\alpha}A_{\beta}\subseteq A_{\alpha+\beta}$ then $A$ can be given the structure of a Lie colour algebra with bracket
[…]
Here, $\varepsilon\colon\mathfrak{g}\times\mathfrak{g}\to\mathbb{F}$ (where $\mathbb{F}$ is the field of scalars) is called a commutation factor [44] (or antisymmetric bicharacter [62], or phase function [46]). That this generalizes a Lie superalgebra is apparent: a Lie superalgebra is simply a Lie colour algebra graded by the group $\Gamma=\mathbb{Z}_{2}$ with commutation factor $\varepsilon(\alpha,\beta)=(-1)^{\alpha\cdot\beta}$. Note that, in general, a commutation factor must satisfy certain conditions (see [44]).
We can define notions such as homomorphisms, subalgebras and representations similarly to superalgebras. However, to retain information about the $\Gamma$-grading, we require some maps (such as homomorphisms, intertwiners, and inclusion maps of subalgebras and $\Gamma$-graded submodules) to be homogeneous of degree 00, as per the following definition.
Definition 2.1.
Let $V=\bigoplus_{\gamma\in\Gamma}V_{\gamma}$ and $W=\bigoplus_{\gamma\in\Gamma}W_{\gamma}$ be vector spaces over $\mathbb{F}$. A linear map $f\colon V\to W$ is homogeneous of degree $\gamma\in\Gamma$ if $f(V_{\alpha})\subseteq W_{\alpha+\gamma}$ for all $\alpha\in\Gamma$. (In particular, $f$ is a homogeneous element of the colour algebra $\mathfrak{gl}(V,\varepsilon)$, see [44].)
2.1. Representations
Although representations can be defined similarly to those for superalgebras, we make an important distinction between graded and ungraded representations.
Definition 2.2.
[…]
A $\Gamma$-graded representation of $\mathfrak{g}$ on a $\Gamma$-graded vector space $V$ is an ungraded representation that also satisfies $\rho(\mathfrak{g}_{\alpha})V_{\gamma}\subseteq V_{\alpha+\gamma}$.
In either case, we say that $V$ is a $\mathfrak{g}$-module given by $\rho$ (or simply a $\mathfrak{g}$-module). We will often omit the representation map $\rho$.
Obviously, we do not require maps related to ungraded representations (such as intertwiners and submodules) to be homogeneous of degree 00. But we do require this condition of the corresponding maps for graded representations.
We call a ($\Gamma$-graded) $\mathfrak{g}$-module ($\Gamma$-graded) irreducible if it has no proper non-trivial ($\Gamma$-graded) submodules. Note that there is an important distinction between the definitions of (ungraded) irreducible and $\Gamma$-graded irreducible: a module may be $\Gamma$-graded irreducible but not ungraded irreducible!
Recall that a $\mathfrak{g}$-module is called completely reducible if it can be written as a direct sum of irreducible submodules. Weyl’s Theorem on complete reducibility will not hold for a general Lie colour algebra (indeed, Weyl’s Theorem does not hold for Lie superalgebras [64]). As such, we need to manually determine when a $\mathfrak{g}$-module is completely reducible. The following propositions will be of great help.
Proposition 2.3.
Consider a ($\Gamma$-graded or ungraded) $\mathfrak{g}$-module $V$ that can be written as $V=\sum_{i\in I}V_{i}$ for irreducible $\mathfrak{g}$-modules $V_{i}$ ($i\in I$). Then, there exists some subset $J\subseteq I$ such that $V=\bigoplus_{i\in J}V_{i}$.
Proposition 2.4.
Submodules and quotient modules of a ($\Gamma$-graded or ungraded) completely reducible $\mathfrak{g}$-module are completely reducible.
These propositions are well-known facts from the module theory of rings (for the proofs in the case of rings, see e.g. [65, Chapter XVII, Lemma 2.1 and Proposition 2.2]). The proofs of these propositions for Lie colour algebra modules are exactly the same as for modules over rings.
2.2. Graded quotient modules
Let $V=\bigoplus_{\gamma\in\Gamma}V_{\gamma}$ be a $\Gamma$-graded $\mathfrak{g}$-module given by the representation $\rho$ and $U=\bigoplus_{\gamma\in\Gamma}U_{\gamma}$ a $\Gamma$-graded submodule of $V$ (so that $U_{\gamma}\subseteq V_{\gamma}$).
Definition 2.5.
The graded quotient module of $V$ by $U$, denoted $V/U$, is a direct sum of quotient vector spaces $\bigoplus_{\gamma\in\Gamma}V_{\gamma}/U_{\gamma}$ with $\mathfrak{g}$-module structure given by $\pi\colon\mathfrak{g}\to\operatorname{End}(V/U)$ defined by
$\pi(x)(v+U_{\gamma})=\rho(x)v+U_{\alpha+\gamma}\qquad\text{for}~{}x\in\mathfrak{g}_{\alpha},\,v\in V_{\gamma}.$
The proof that the above representation is well defined is similar to that for Lie superalgebras. The only additional property to check is whether $\pi$ is homogeneous of degree 00, which is easily verified.
The three isomorphism theorems hold for $\mathfrak{g}$-modules. Again, the proofs are similar to superalgebras, needing only to verify that all maps are homogeneous of degree 00.
2.3. Discolouration
Scheunert [44] gave a bijection between Lie colour algebras and graded Lie superalgebras and between their representations. This bijection has many nice properties; for example, it preserves subalgebras, ideals and subrepresentations. To compute the image of a colour algebra under this bijection, we deform the bracket using a multiplier.
[…]
Note that there is an alternative method of discolouration which works by adding new elements, called Klein operators, to the universal enveloping algebra (see [45, 46, 47, 48]). However, for the remainder of this paper we will use the original discolouration process of [44] described in this section.
It is also possible to discolour representations. Given a graded representation $\rho\colon\mathfrak{g}\to\mathfrak{gl}(V,\varepsilon)$ and a discolouring multiplier $\sigma\colon\Gamma\times\Gamma\to\mathbb{F}^{\times}$, we can define the discolouration of $\rho$ to be the graded representation $\rho^{\sigma}$ of $\mathfrak{g}^{\sigma}$ given by
$\rho^{\sigma}(x)v=\sigma(\alpha,\gamma)\rho(x)v,\qquad\text{for}~{}x\in\mathfrak{g}_{\alpha},v\in V_{\gamma}.$
Similarly, the map $\rho\mapsto\rho^{\sigma}$ is a bijection between the graded representations of $\mathfrak{g}$ and the graded representations of the corresponding discoloured $\Gamma$-graded Lie superalgebra [44]. Thus, the representation theory of Lie colour algebras can be completely derived from that of $\Gamma$-graded Lie superalgebras.
The $\mathbb{Z}_{2}$-graded irreducible representations of Lie superalgebras have been well studied. However, these results cannot be immediately applied to find $\Gamma$-graded irreducible representations. For instance, it might not be possible to find a $\Gamma$-grading for a $\mathbb{Z}_{2}$-graded representation. On the other hand, there might be a $\Gamma$-graded irreducible representation that becomes reducible when considered as a $\mathbb{Z}_{2}$-graded representation.
3. loop module
Throughout the remainder of this paper, we will assume that $\Gamma$ is a finite additive abelian group and that the ambient field $\mathbb{F}$ contains a primitive root of unity of order |$\absolutevalue{\Gamma}$ | (for instance, $\mathbb{F}=\mathbb{C}$ will do).
Let $\mathfrak{g}=\bigoplus_{\gamma\in\Gamma}\mathfrak{g}_{\gamma}$ be a $\Gamma$-graded Lie colour algebra with commutation factor $\varepsilon$ and let $H$ be a subgroup of the (additive) abelian group $\Gamma$. Notice that $\mathfrak{g}$ has a natural $\Gamma/H$ grading, given by $\mathfrak{g}=\bigoplus_{\Lambda\in\Gamma/H}\mathfrak{g}^{\Gamma/H}_{\Lambda}$ where $\mathfrak{g}^{\Gamma/H}_{\Lambda}=\bigoplus_{\alpha\in\Lambda}\mathfrak{g}_{\alpha}$. Recall [44] that $\mathfrak{g}$ also has a natural $\mathbb{Z}_{2}$-grading
$\mathfrak{g}^{0}=\bigoplus_{\gamma\in\Gamma_{0}}\mathfrak{g}_{\gamma},\qquad\mathfrak{g}^{1}=\bigoplus_{\gamma\in\Gamma_{1}}\mathfrak{g}_{\gamma}.$
where $\Gamma_{0}=\{\gamma\in\Gamma\mid\varepsilon(\gamma,\gamma)=1\}$ and $\Gamma_{1}=\Gamma\setminus\Gamma_{0}$. If we choose $H\leq\Gamma_{0}$, then the $\Gamma/H$-grading refines the $\mathbb{Z}_{2}$-grading, i.e.
$\mathfrak{g}^{0}=\bigoplus_{\Lambda\in\Gamma_{0}/H}\mathfrak{g}^{\Gamma/H}_{\Lambda},\qquad\mathfrak{g}^{1}=\bigoplus_{\Lambda\in\Gamma_{1}/H}\mathfrak{g}^{\Gamma/H}_{\Lambda}.$
This is useful if we want to prevent $H$ from interfering with discolouration.
Definition 3.1 (Mazorchuk–Zhao [58]; see also [60]).
Let $V=\bigoplus_{\Lambda\in\Gamma/H}V_{\Lambda}$ be a $\Gamma/H$-graded $\mathfrak{g}$-module given by $\rho$. Recall that the group algebra $\mathbb{F}\Gamma$ is the algebra with basis $\{e_{\gamma}\mid\gamma\in\Gamma\}$ and algebra product defined by $e_{\alpha}\cdot e_{\beta}=e_{\alpha+\beta}$. The vector space $V\otimes\mathbb{F}\Gamma$ can be given a $\mathfrak{g}$-module structure by
$x(v\otimes e_{\beta})=xv\otimes e_{\alpha+\beta}$
for $x\in\mathfrak{g}_{\alpha}$ and $v\otimes e_{\beta}\in V\otimes\mathbb{F}\Gamma$. The Loop module of $V$ by $H$ is the $\Gamma$-graded $\mathfrak{g}$-module
$\mathrm{L}_{H}{(V)}=\bigoplus_{\gamma\in\Gamma}V_{\gamma+H}\otimes e_{\gamma}\subset V\otimes\mathbb{F}\Gamma.$
The $\gamma$-sector of $\mathrm{L}_{H}{(V)}$ is $\mathrm{L}_{H}{(V)}_{\gamma}=V_{\gamma+H}\otimes e_{\gamma}$.
Not every $\Gamma/H$-graded module $V$ can be given a $\Gamma$-grading. Intuitively, the loop module adds in multiple copies of $V$ to ensure that there is enough ‘space’ for a refined $\Gamma$-grading. If $V$ can be given a $\Gamma$-grading, then the loop module contains $V$:
Proposition 3.2.
If $V$ can be given a $\Gamma$-grading that refines the $\Gamma/H$-grading, then $V$ is a $\Gamma$-graded submodule of $\mathrm{L}_{H}{(V)}$.
Proof.
Let $V=\bigoplus_{\gamma\in\Gamma}W_{\gamma}$ be the $\Gamma$-grading, with $V_{\Lambda}=\bigoplus_{\gamma\in\Lambda}W_{\gamma}$ for all $\Lambda\in\Gamma/H$ (i.e. the $\Gamma$-grading refines the $\Gamma/H$ grading). Define a map $\Phi\colon V\to\mathrm{L}_{H}{(V)}$ by $\Phi(w)=w\otimes e_{\gamma}$ for $w\in W_{\gamma}$ and extending to inhomogeneous elements by linearity. Linearity and injectivity of $\Phi$ follow immediately. The map $\Phi$ is homogeneous of degree 00 by definition. And $\Phi$ is an intertwiner: for $w\in W_{\gamma}$ and $x\in\mathfrak{g}_{\alpha}$,
$\Phi(xw)=xw\otimes e_{\alpha+\gamma}=x(w\otimes e_{\gamma})=x(\Phi(w))\qed$
As well as a $\Gamma$-grading, $\mathrm{L}_{H}{(V)}$ has a natural $\Gamma/H$-grading given by
$\mathrm{L}_{H}{(V)}=\bigoplus_{\Lambda\in\Gamma/H}\left(\bigoplus_{\gamma\in\Lambda}V_{\Lambda}\otimes e_{\gamma}\right).$
Remark 3.3.
If $\Gamma=K\times H$ for some group $K$, then we could equivalently define the loop module to be $V\otimes\mathbb{F}[H]$ with representation given by $x(v\otimes e_{h})=xv\otimes(e_{h+\eta})$ for $x\in\mathfrak{g}_{(\kappa,\eta)}$ and $v\otimes e_{h}\in V\otimes\mathbb{F}[H]$. Compare with [59], where this construction, together with a twist by a multiplier $\sigma$ on $\mathbb{F}[H]$, is called an enhancement. The enhancement approach is used in [36] and [59] to study the structure of simple Lie colour algebras themselves (rather than their modules). An enhancement of a Lie colour algebra is similar to the loop algebra construction studied elsewhere [60, 58].
3.1. loop module and quantum mechanical systems
A common way to obtain a $\Gamma$-graded colour supersymmetric quantum mechanical system is by modifying an existing supersymmetric system. This is the approach taken in (for example) [23], where the $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$-graded supersymmetric system presented in [23, Section III] can be obtained by introducing an additional central charge to Witten’s model [66, Section 6]. In [23], the Hilbert space realisation for this $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$-graded system was constructed independently from that of Witten’s model. In this section, we will demonstrate that this $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$-graded Hilbert space realisation can be constructed from Witten’s realisation using a loop module. In this way, the loop module has already appeared in the Lie colour algebra literature and is a natural structure to consider in the study of Lie colour algebras.
Witten considered [66, Section 6] a superalgebra with even sector spanned by a Hamiltonian $H$ and odd sector spanned by two supercharges $Q_{1}$ and $Q_{2}$ satisfying
[…]
for $i,j=1,2$. This algebra was realised as operators acting on two component Pauli spinors in the Hilbert space $\mathscr{H}^{\textup{W}}=L^{2}(\mathbb{R})\otimes\mathbb{C}^{2}$ via
[…]
By adding a central charge,
[…]
the original superalgebra becomes a $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$-graded colour Lie superalgebra (without requiring recolouration) with sectors spanned by
[…]
and satisfying the following additional relations
[…]
while all other Lie colour brackets vanish. We can easily verify that $\mathscr{H}^{\textup{W}}$ becomes an ungraded module of the Lie colour algebra.
Taking the loop module $\mathrm{L}_{\mathbb{Z}_{2}}{(\mathscr{H}^{\textup{W}})}$ turns $\mathscr{H}^{\textup{W}}$ into a $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$-graded module. Here, we use $\mathbb{Z}_{2}\cong\{00,11\}\leq\mathbb{Z}_{2}\times\mathbb{Z}_{2}$. Set
$\mathscr{H}_{00}=\mathscr{H}^{\textup{W}}_{0}\otimes e_{00}\qquad\mathscr{H}_{11}=\mathscr{H}^{\textup{W}}_{0}\otimes e_{11}\qquad\mathscr{H}_{01}=\mathscr{H}^{\textup{W}}_{1}\otimes e_{01}\qquad\mathscr{H}_{10}=\mathscr{H}^{\textup{W}}_{1}\otimes e_{10}.$
Then, the loop module is $\mathrm{L}_{\mathbb{Z}_{2}}{(\mathscr{H}^{\textup{W}})}=\mathscr{H}_{00}\oplus\mathscr{H}_{01}\oplus\mathscr{H}_{11}\oplus\mathscr{H}_{10}=L^{2}(\mathbb{R})\otimes\mathbb{C}^{4}$ (note the order of the sectors). We can then explicitly compute the representation on four-component spinors:
$\rho^{\mathbb{Z}_{2}}(H)$ $=\begin{pmatrix}H&0\\ 0&H\end{pmatrix},$ $\rho^{\mathbb{Z}_{2}}(Q_{1})$ $=\begin{pmatrix}Q_{1}&0\\ 0&Q_{1}\end{pmatrix},$ $\rho^{\mathbb{Z}_{2}}(Q_{2})$ $=\begin{pmatrix}0&Q_{2}\\ Q_{2}&0\end{pmatrix},$ $\rho^{\mathbb{Z}_{2}}(Z)$ $=\begin{pmatrix}0&Z\\ Z&0\end{pmatrix}$
where $\rho^{\mathbb{Z}_{2}}$ is the representation for $\mathrm{L}_{\mathbb{Z}_{2}}{(\mathscr{H}^{\textup{W}})}$. Setting $H_{00}=2\rho^{\mathbb{Z}_{2}}(H),\,Q_{01}=\rho^{\mathbb{Z}_{2}}(Q_{1}),\,Q_{10}=\rho^{\mathbb{Z}_{2}}(Q_{2})$ and $Z_{11}=2\rho^{\mathbb{Z}_{2}}(Z)$, we obtain the $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$-supersymmetry model in [23, Section III] (with $m=1/2$).
The above example demonstrates that the loop module provides a natural method for constructing the Hilbert space of a $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$-graded supersymmetry model from a $\mathbb{Z}_{2}$-graded supersymmetry model. The loop module can be applied to other colour supersymmetric models constructed in a similar way.
3.2. Loop module and irreducibility
Under the assumptions outlined at the beginning of this section, and for $\mathfrak{g}$ a $\Gamma$-graded Lie algebra (i.e. with trivial commutation factor $\varepsilon(\alpha,\beta)=1$ for all $\alpha,\beta\in\Gamma$), [58] showed that every $\Gamma$-graded irreducible module can be obtained from the non-graded ones via the loop module:
Theorem 3.4 (Mazorchuk–Zhao [58, Theorem 24]).
Let $W$ be a $\Gamma$-graded irreducible $\mathfrak{g}$-module. There exists a subgroup $H\leq\Gamma$ and an ungraded irreducible $\mathfrak{g}$-module $V$ such that $W$ is isomorphic to $\mathrm{L}_{H}{(V)}$ as a $\Gamma$-graded module.
Note that this theorem still holds if $\mathfrak{g}$ is instead a Lie colour algebra with non-trivial commutation factor (e.g. if $\mathfrak{g}$ is a $\mathbb{Z}_{2}$-graded Lie superalgebra). Indeed, the proof works for modules over any associative algebra; in particular, over the universal enveloping algebra for $\mathfrak{g}$. A similar result [58, Theorem 31] is available if $\Gamma$ is an infinite group, but we will restrict our attention to the finite case in this paper.
The above theorem reduces the study of $\Gamma$-graded irreducible modules to the study of ungraded modules. In combination with discolouration, this allows us to derive all such $\Gamma$-graded irreducible colour algebra representations from the $\mathbb{Z}_{2}$-graded irreducible superalgebra representations.
However, one difficulty with applying Theorem 3.4 in practice is that the choice of $H$ in the theorem is not necessarily unique i.e. potentially $W\cong\mathrm{L}_{H_{1}}{(V)}\cong\mathrm{L}_{H_{2}}{(V)}$ for non-isomorphic subgroups $H_{1},H_{2}\leq G$. Determining when two loop modules are isomorphic is an open problem [58, Problem 36].
Building on the work of [58], it was shown that the loop module construction defines a functor and can be extended to an equivalence between two categories (see [61] for details). Moreover, using Brauer invariants, it was shown that the loop module defines a bijection between equivalence classes of $\Gamma/H$-graded irreducible and $\Gamma$-graded irreducible modules [61, Remark 7.2].
4. Iterated loop module bijection
In this section, we will present a bijection between equivalence classes of $\Gamma/H$- and $\Gamma$-graded irreducible representations, similar to that of [61]. However, instead of using Brauer invariants, we will provide an alternate iterative approach. Given a finite grading group $\Gamma$, we suggest using the Jordan–Hölder decomposition
(1) $\Gamma\rhd N_{\ell}\rhd N_{\ell-1}\rhd\cdots\rhd N_{1}\rhd\{0\}$
and looking at submodules of the following iterated loop module
(2) $\mathrm{L}_{N_{1}/\{0\}}{(\cdots\mathrm{L}_{N_{\ell}/N_{\ell-1}}{(\mathrm{L}_{\Gamma/N_{\ell}}{(V)})}\cdots)}.$
(Note that we are making use of the isomorphism $\Gamma/N_{\ell}\cong(\Gamma/N_{\ell-1})/(N_{\ell}/N_{\ell-1})$ (for instance)). Note that this iterative approach uses only simple modules, which greatly simplifies the possible options.
Our strategy is as follows: we first prove that every finite-dimensional $\Gamma$-graded irreducible module appears as a quotient of a loop module (Lemma 4.1). The loop module is completely reducible both as a $\Gamma/H$- and $\Gamma$-graded module (Lemmas 4.2 and 4.5 respectively) and we explicitly provide the direct sum decompositions. Comparing these direct sum decompositions limits the possible $\Gamma$-graded irreducible modules to either $\Gamma/H$-graded irreducible modules or their loop modules. We use these results to construct the bijection between $\Gamma/H$- and $\Gamma$-graded modules (Theorem 4.8) up to equivalence defined by the irreducible direct summands.
4.1. Quotient modules of the loop module
As in Section 3, let $\mathfrak{g}$ be a $\Gamma$-graded Lie colour algebra and $V$ be a finite-dimensional $\Gamma/H$-graded irreducible $\mathfrak{g}$-module. With the iterated loop module construction of (2) in mind, we will assume that $H$ is simple. Note that, since $H$ is abelian, finite and simple, we must have that $H\cong\mathbb{Z}_{p}$ for some prime $p$.
In general, the loop module $\mathrm{L}_{H}{(V)}$ will not be $\Gamma$-graded irreducible. However, to construct such a $\Gamma$-graded irreducible $\mathfrak{g}$-module, we can simply quotient out by a $\Gamma$-graded maximal submodule. Since $\mathrm{L}_{H}{(V)}$ may have many $\Gamma$-graded maximal submodules, this procedure may yield many nonisomorphic $\Gamma$-graded modules for each $\Gamma/H$-graded maximal submodule. Regardless, we will show that every $\Gamma$-graded irreducible $\mathfrak{g}$-module can be constructed in this way.
Lemma 4.1.
Let $\mathfrak{g}=\bigoplus_{\gamma\in\Gamma}\mathfrak{g}_{\gamma}$ be a Lie colour algebra. Let $W=\bigoplus_{\gamma\in\Gamma}W_{\gamma}$ be a finite-dimensional $\Gamma$-graded irreducible $\mathfrak{g}$-module. Then, for $H\leq\Gamma$, there exists a $\Gamma/H$-graded irreducible $\mathfrak{g}$-module $V$ such that $W$ is a graded quotient of $\mathrm{L}_{H}{(V)}$.
Proof.
Note that $W$ has a natural $\Gamma/H$-grading given by
$\bigoplus_{\Lambda\in\Gamma/H}\left(\bigoplus_{\gamma\in\Lambda}W_{\gamma}\right).$
Let $V$ be a $\Gamma/H$-graded irreducible submodule of $W$. Such a $V$ exists by strong induction on the dimension of $W$. (See also [58, Lemma 23].)
Take $v\in V_{\Lambda}$ and write
(3) $v=\sum_{\gamma\in\Lambda}w_{\gamma}$
for $w_{\gamma}\in W_{\gamma}$. For $\gamma\in\Lambda$, let $\pi_{\gamma}\colon V_{\Lambda}\to W_{\gamma}$ be the projection of $V_{\Lambda}$ onto $W_{\gamma}$,
$\pi_{\gamma}(v)=w_{\gamma}.$
Note that $\pi_{\gamma}$ is well-defined because the linear combination in (3) is unique due to the direct sum $V_{\Lambda}\subseteq\bigoplus_{\gamma\in\Lambda}W_{\gamma}$. Additionally, it is easy to see that $\pi_{\gamma}$ is linear.
Note that, for $x\in\mathfrak{g}_{\alpha}$, we have $xv=\sum_{\gamma\in\Lambda}x\pi_{\gamma}(v)$. Since $x\pi_{\gamma}(v)\in W_{\alpha+\gamma}$ and such a linear combination of homogeneous elements is unique, we have that $\pi_{\alpha+\gamma}(xv)=x\pi_{\gamma}(v)$ by the definition of $\pi_{\alpha+\gamma}$.
Define a map $\Phi\colon\mathrm{L}_{H}{(V)}\to W$ by
$\Phi(v\otimes e_{\gamma})=\pi_{\gamma}(v).$
Then
$x\Phi(v\otimes e_{\gamma})=x\pi_{\gamma}(v)=\pi_{\alpha+\gamma}(xv)=\Phi(x(v\otimes e_{\gamma}))$
so $\Phi$ is in intertwiner.
Therefore, by the First Isomorphism Theorem, $\operatorname{im}\Phi\cong\mathrm{L}_{H}{(V)}/\ker\Phi$. But $\operatorname{im}\Phi$ is a $\Gamma$-graded submodule of $W$, so $\operatorname{im}\Phi=0$ or $\operatorname{im}\Phi=W$ by irreducibility of $W$. However, since $V\neq 0$, and $\Phi$ is defined using projections, we clearly have $\operatorname{im}\Phi\neq 0$. We conclude that $\operatorname{im}\Phi=W$; hence $W\cong\mathrm{L}_{H}{(V)}/\ker\Phi$ as claimed. ∎
Lemma 4.1 provides us with a general method of obtaining the $\Gamma$-graded irreducible modules from the $\Gamma/H$-graded ones; namely find all $\Gamma$-graded maximal submodules of $\mathrm{L}_{H}{(V)}$ and take quotients. Note that we can first fix $H$ and then find all $\Gamma$-graded irreducible modules. Compare this with Theorem 3.4, where we must choose a different $H$ for each $\Gamma$-graded module.
4.2. $\Gamma/H$-grading: loop module is completely reducible
The loop module is a completely reducible $\Gamma/H$-graded module. It is shown in [58] that the irreducible direct summands are isomorphic to $V$ as vector spaces, but the $\mathfrak{g}$-action is twisted by an element of the character group $\widehat{\Gamma}$. These direct summands will restrict which quotients are possible, and hence which $\Gamma$-graded irreducible representations are possible in Lemma 4.1. We will now briefly recount these results.
Let $\widehat{\Gamma}$ be the character group: the group of all homomorphisms $\widehat{\Gamma}\to\mathbb{F}^{\times}$ with pointwise multiplication the group operation. Our earlier assumption that $\mathbb{F}$ contains a primitive root of unity of order |$\absolutevalue{\Gamma}$ | ensures that $\widehat{\Gamma}\cong\Gamma$. Let $H^{\perp}=\{f\in\widehat{\Gamma}\mid f(\eta)=1~{}\text{for all}~{}\eta\in H\}$. Note that since $\Gamma$ is finite, |$\absolutevalue{\widehat{\Gamma}/H^{\perp}}=\absolutevalue{H}$ |.
Let $V=\bigoplus_{\Lambda\in\Gamma/H}V_{\Lambda}$ be a $\Gamma/H$-graded $\mathfrak{g}$-module with representation $\rho\colon\mathfrak{g}\to\operatorname{End}(V)$. For each $fH^{\perp}\in\widehat{\Gamma}/H^{\perp}$, we define the twisted module $V^{f}$ to be equal to $V$ as a vector space, but with the following twisted representation $\rho^{f}\colon\mathfrak{g}\to\operatorname{End}(V^{f})$
$\rho^{f}(x)v=f(\alpha)\rho(x)v\qquad\text{for}~{}x\in\mathfrak{g}_{\alpha},\,v\in V.$
Note that $f(\alpha)$ will be a root of unity.
As a $\Gamma/H$-module, the loop module is a direct sum of all such twisted modules $V^{f}$:
Lemma 4.2 (Mazorchuk–Zhao [58, Lemma 20]).
Let $H$ be a finite group and $V=\bigoplus_{\Lambda\in\Gamma/H}V_{\Lambda}$ be a $\Gamma/H$-graded module of a Lie colour algebra $\mathfrak{g}$ over $\mathbb{F}$. Assume that $\mathbb{F}$ has a primitive root of unity of order |$\absolutevalue{\Gamma}$ |. As a $\Gamma/H$-graded module, the loop module $\mathrm{L}_{H}{(V)}$ is isomorphic to
$\bigoplus_{f\in\widehat{\Gamma}/H^{\perp}}V^{f}.$
If we assume that $V$ is irreducible, then Lemma 4.2 (together with 4.3 (i) below) shows that $\mathrm{L}_{H}{(V)}$ is completely reducible as a $\Gamma/H$-graded module and gives a direct sum decomposition. In combination with Lemma 4.1, this restricts the possible structures of the $\Gamma$-graded irreducible modules.
For use in later proofs, we now prove three important properties of $V^{f}$.
Proposition 4.3.
The following three properties hold:
  1. (i)
    If $V$ is $\Gamma/H$-graded irreducible then $V^{f}$ is $\Gamma/H$-graded irreducible.
  2. (ii)
    There is an equality of $\mathfrak{g}$-modules: $(\mathrm{L}_{H}{(V)})^{f}=\mathrm{L}_{H}{(V^{f})}$.
  3. (iii)
    If $V$ can be given a $\Gamma$-grading that refines the $\Gamma/H$-grading, then $V^{f}\cong V$.
Proof.
Since the action of $\rho$ only differs from $\rho^{f}$ by a scalar (at most), $\rho(\mathfrak{g})U\subseteq U$ if and only if $\rho^{f}(\mathfrak{g})U\subset U$ for any submodule $U$ of $V$ or $V^{f}$. Property (i) follows.
For (ii), $(\mathrm{L}_{H}{(V)})^{f}$ equals $\mathrm{L}_{H}{(V^{f})}$ as a vector space and both of these modules have the same representation: for $x\in\mathfrak{g}_{\alpha}$ and $v\in V_{\gamma+H}$,
$x\cdot(v\otimes e_{\gamma})=f(\alpha)\rho(x)v\otimes e_{\gamma}.$
To show (iii), let $V=\bigoplus_{\gamma\in\Gamma}W_{\gamma}$ be the $\Gamma$-grading, with $V_{\Lambda}=\bigoplus_{\gamma\in\Lambda}W_{\gamma}$ for all $\Lambda\in\Gamma/H$ (i.e. the $\Gamma$-grading refines the $\Gamma/H$ grading). Take $v\in W_{\gamma}$ and define $\Phi\colon V\to V^{f}$ by
$\Phi(v)=f(\gamma)v.$
Obviously, $\Phi$ is a linear isomorphism and homogeneous of degree 00. Additionally, $\Phi$ is an intertwiner:
$\Phi(\rho(x)v)=f(\alpha+\gamma)\rho(x)v=f(\alpha)\rho(x)f(\gamma)v=\rho^{f}(x)\Phi(v)$
completing the proof. ∎
It is clear from the definition and the above propositions that $V^{f}$ has a very similar module structure to $V$. This will be useful when examining the structure of $\mathrm{L}_{H}{(V)}$.
4.3. $\Gamma$-grading: loop module is completely reducible
We will now show that the loop module is completely reducible as a $\Gamma$-graded module. Similar to $\Gamma/H$-graded case, the $\Gamma$-graded irreducible direct summands all have a similar structure: they are all (what we will call) $H$-parity shifts of the same $\Gamma$-graded irreducible module. A $H$-parity shift is obtained by simply applying a circular shift to each of the sectors.
Comparing the $\Gamma$- and $\Gamma/H$-graded irreducible direct summands will give us more information about the structure of the loop module. This will allow us to conclude that the only possible $\Gamma$-graded direct summands of $\mathrm{L}_{H}{(V)}$ are $V$ with a $\Gamma$-grading or $\mathrm{L}_{H}{(V)}$ itself.
4.3.1. Direct sum decomposition
Take some $h\in H$ and a $\Gamma$-graded module $W$. We can define the $H$-parity shift of $W$ by $h$ as a new $\Gamma$-graded module $W^{+h}=\bigoplus_{\gamma\in\Gamma}W^{+h}_{\gamma}$, where $W^{+h}_{\gamma}=W_{\gamma+h}$. Obviously, a $H$-parity shift of $W$ is still a $\Gamma$-graded $\mathfrak{g}$-module.
Following this definition, we make three immediate observations:
  1. (1)
    A $H$-parity shift might yield a module isomorphic to the original—for example, any $H$-parity shift of the loop module $\mathrm{L}_{H}{(V)}$ is isomorphic to $\mathrm{L}_{H}{(V)}$.
  2. (2)
    If $W$ is irreducible then $W^{+h}$ is irreducible.
  3. (3)
    The $\mathfrak{g}$-modules $W$ and $W^{+h}$ are isomorphic as $\Gamma/H$-graded modules.
Proposition 4.4.
Let $V$ be a $\Gamma/H$-graded module and $W$ be a $\Gamma$-graded submodule of $\mathrm{L}_{H}{(V)}$. Then $W^{+h}$ can be embedded in $\mathrm{L}_{H}{(V)}$.
Proof.
Define a map $\Psi\colon W^{+h}\to\mathrm{L}_{H}{(V)}$ by
$\Psi(w\otimes e_{\gamma+h})=w\otimes e_{\gamma}$
for $w\otimes e_{\gamma+h}\in W^{+h}_{\gamma}$ and extending to inhomogeneous elements by linearity. This map is obviously an embedding of $W^{+h}$ into $\mathrm{L}_{H}{(V)}$. ∎
From now on, we will identify $W^{+h}$ with its embedding in $\mathrm{L}_{H}{(V)}$.
Lemma 4.5.
Let $H\leq\Gamma$ be a finite simple group and $V=\bigoplus_{\Lambda\in\Gamma/H}V_{\Lambda}$ be a finite-dimensional $\Gamma/H$-graded irreducible module of a Lie colour algebra $\mathfrak{g}$ over $\mathbb{F}$. Then, the loop module $\mathrm{L}_{H}{(V)}$ is completely reducible as a $\Gamma$-graded module with
$\mathrm{L}_{H}{(V)}=\bigoplus_{h\in K}W^{+h}$
for any $\Gamma$-graded irreducible module $W\subseteq\mathrm{L}_{H}{(V)}$ and some subset $K\subseteq H$.
Proof.
Take any $\Gamma$-graded irreducible module $W\subseteq\mathrm{L}_{H}{(V)}$. Such a module exists because $\mathrm{L}_{H}{(V)}$ is finite-dimensional.
Define a map $\Phi\colon W\to V$ by
$\Phi(w\otimes e_{\gamma})=w$
for $w\otimes e_{\gamma}\in W_{\gamma}\subseteq\mathrm{L}_{H}{(V)}_{\gamma}$ and extending to inhomogeneous elements by linearity. We claim that $\Phi$ is a $\Gamma/H$-graded intertwiner. That $\Phi$ is linear and homogeneous of degree 00 is obvious. For $x\in\mathfrak{g}_{\alpha}$, we have
$\Phi(x(w\otimes e_{\gamma}))=\Phi(xw\otimes e_{\alpha+\gamma})=xw=x\Phi(w\otimes e_{\gamma})$
which proves the claim.
We now know that $\Phi(W)$ is a $\Gamma/H$-graded submodule of $V$. But since $W\neq 0$, we must have that $\Phi(W)\neq 0$. Since $V$ is $\Gamma/H$-graded irreducible, we must have that $\Phi(W)=V$.
Now, consider the module $\Sigma_{h\in H}W^{+h}\subseteq\mathrm{L}_{H}{(V)}$. In particular, consider the $\gamma$-sector $\sum_{h\in H}W^{+h}_{\gamma}$ for some $\gamma\in\Gamma$. Applying the map $\Phi$,
$\Phi\left(\sum_{h\in H}W^{+h}_{\gamma}\right)=\sum_{h\in H}\Phi\left(W^{+h}_{\gamma}\right)=\sum_{h\in H}\Phi\left(W_{\gamma+h}\right)=\Phi\left(\bigoplus_{h\in H}W_{\gamma+h}\right)=V_{\gamma+H}$
using the fact $\Phi(W)=V$ (and recalling the embedding in 4.4). By examining the definition of $\Phi$, we find tha $\sum_{h\in H}W^{+h}_{\gamma}=V_{\gamma+H}\otimes e_{\gamma}=\mathrm{L}_{H}{(V)}_{\gamma}$. Therefore, $\sum_{h\in H}W^{+h}=\mathrm{L}_{H}{(V)}$.
All that remains to show is that we can take a subset of $H$ to obtain a direct sum. But this follows from 2.3. ∎
Theorem 4.6.
Let $V$ be $\Gamma/H$-graded irreducible. Every $\Gamma$-graded irreducible submodule of $\mathrm{L}_{H}{(V)}$ is isomorphic as a $\Gamma/H$-graded submodule to either $\mathrm{L}_{H}{(V)}$ or $V$.
Proof.
Take $W$ a $\Gamma$-graded irreducible submodule of $\mathrm{L}_{H}{(V)}$. By Lemma 4.5,
$\mathrm{L}_{H}{(V)}=\bigoplus_{h\in K}W^{+h}.$
Since $\mathrm{L}_{H}{(V)}$ is a completely reducible $\Gamma/H$-graded module, we know that $W$ is a completely reducible $\Gamma/H$-graded module with direct summands contained in those of $\mathrm{L}_{H}{(V)}$. By Lemma 4.2, $W$ is a direct sum of character twists $V^{f}$ ($f\in\widehat{\Gamma}/H^{\perp}$). Let $n$ be the number of $\Gamma/H$-graded direct summands in $W$.
[…]
In combination with Lemma 4.1, the above theorem tells us that every $\Gamma$-graded module appears as either a $\Gamma/H$-graded irreducible module or the loop module of one. This result will greatly aid in the search for graded irreducible Lie colour algebra representations. Noting that $V$ is isomorphic to $\mathrm{L}_{\{0\}}{(V)}$, Theorem 4.6 is a similar result as [58, Theorme 24] (restated in Theorem 3.4): every $\Gamma$-graded module is a loop module. However, there are two important distinctions between Theorem 4.6 and [58, Theorem 24]: restricting $H$ to a simple group greatly restricts the possible loop modules, and $H$ does not rely on the choice of $V$.
One difficulty remains: in the former case, finding a $\Gamma$-grade for a $\Gamma/H$-graded module can be difficult in general. For some results about the gradability of modules see [56]. To aid with this, we prove a uniqueness result for the $\Gamma$-grade.
4.3.2. Uniqueness of the $\Gamma$-grade
It is possible for two modules to be non-isomorphic as $\Gamma$-graded modules but isomorphic as $\Gamma/H$-graded modules. This is important to keep in mind when trying to find a $\Gamma$-grading for a $\Gamma/H$-graded irreducible module; there might be multiple different $\Gamma$-gradings, all of which must be considered. However, the following proposition gives a uniqueness result for the $\Gamma$-grade.
Proposition 4.7.
If $V$ is a finite-dimensional $\Gamma/H$-graded irreducible $\mathfrak{g}$-module, then any two $\Gamma$-gradings for $V$ are equivalent up to $H$-parity shift.
More precisely, this proposition tells us that if $V=\bigoplus_{\gamma\in\Gamma}V_{\gamma}$ and $V=\bigoplus_{\gamma\in\Gamma}\widetilde{V}_{\gamma}$ are two $\Gamma$-gradings for $V$ which both refine the $\Gamma/H$ grading (i.e. for each $\Lambda\in\Gamma/H$ the original $\Lambda$-sector is $\bigoplus_{\gamma\in\Lambda}V_{\gamma}=\bigoplus_{\gamma\in\Lambda}\widetilde{V}_{\gamma}$) then there exists $\eta\in H$ such that $V^{+\eta}_{\gamma}=\widetilde{V}_{\gamma}$.
Proof.
Both $V=\bigoplus_{\gamma\in\Gamma}V_{\gamma}$ and $V=\bigoplus_{\gamma\in\Gamma}\widetilde{V}_{\gamma}$ appear as $\Gamma$-graded irreducible submodules of $\mathrm{L}_{H}{(V)}$ by 3.2. Since all graded $\Gamma$-graded irreducible direct summands of $\mathrm{L}_{H}{(V)}$ are $H$-parity shifts of $V$ by Lemma 4.5, the result follows. ∎
4.4. The bijection
Let $\mathfrak{g}=\bigoplus_{\gamma\in\Gamma}\mathfrak{g}_{\gamma}$ be a Lie colour algebra and $H\leq\Gamma$ be a finite simple group. It is easily verified that twists by an element of $\widehat{\Gamma}/H^{\perp}$ and $H$-parity shifts both define equivalence relations. Let $\mathsf{Irr}_{\Gamma/H}^{\textup{tw}}(\mathfrak{g})$ be the collection of equivalence classes of finite-dimensional $\Gamma/H$-graded irreducible $\mathfrak{g}$-modules, where two $\Gamma/H$-graded modules are equivalent up to a twist by a character in $\widehat{\Gamma}/H^{\perp}$. Let $\mathsf{Irr}_{\Gamma}^{H\textup{p}}(\mathfrak{g})$ be the collection of equivalence classes of finite-dimensional $\Gamma$-graded irreducible $\mathfrak{g}$-modules, where two $\Gamma$-graded modules are equivalent up to $H$-parity shift.
Theorem 4.8.
Define a function $\mathcal{F}\colon\mathsf{Irr}_{\Gamma/H}^{\textup{tw}}\to\mathsf{Irr}_{\Gamma}^{H\textup{p}}$ by
(4) $\mathcal{F}(V)=\begin{cases}V&\text{if $V$ is $\Gamma$-gradable}\\ \mathrm{L}_{H}{(V)}&\text{otherwise}\end{cases}$
where $V$ and $\mathrm{L}_{H}{(V)}$ are representatives of the equivalence classes. Then $\mathcal{F}$ is well-defined (does not depend on the choice of representative $V$) and bijective.
Proof.
The $\Gamma$-gradability remains unchanged after applying an isomorphism or after twisting by $f\in\widehat{\Gamma}/H^{\perp}$. And any $\Gamma$-grading of $V$ is unique up to $H$-parity shift (4.7). Observe that $\mathcal{F}(V^{f})=(\mathcal{F}(V))^{f}\cong\mathcal{F}(V)$ (using 4.3 (ii)), since the image of $\mathcal{F}$ is $\Gamma$-graded in either case and all twists by $\widehat{\Gamma}/H^{\perp}$ of a $\Gamma$-graded module are isomorphic (4.3 (iii)). We can hence deduce that $\mathcal{F}$ is well defined.
For injectivity, assume that $\mathcal{F}(V)=\mathcal{F}(\widetilde{V})$. Using Lemma 4.2, we know that $V$ is not $\Gamma/H$-graded isomorphic to $\mathrm{L}_{H}{(\widetilde{V})}$, so $V$ cannot be equivalent to $\mathrm{L}_{H}{(\widetilde{V})}$ in $\mathsf{Irr}_{\Gamma}^{H\textup{p}}$ (since a $H$-parity shift is a $\Gamma/H$-graded isomorphism). Thus, if $\mathcal{F}(V)=V$ then $\mathcal{F}(\widetilde{V})=\widetilde{V}$. Similarly, if $\mathcal{F}(\widetilde{V})=\widetilde{V}$ then $\mathcal{F}(V)=V$. But if $\mathcal{F}(V)=V$ and $\mathcal{F}(\widetilde{V})=\widetilde{V}$ are equivalent in $\mathsf{Irr}_{\Gamma}^{H\textup{p}}$, then we must have $V\cong\widetilde{V}$ as $\Gamma/H$-modules. That is, $\mathcal{F}$ is injective in this case
There is one remaining case to check for injectivity: $\mathcal{F}(V)=\mathrm{L}_{H}{(V)}$ and $\mathcal{F}(\widetilde{V})=\mathrm{L}_{H}{(\widetilde{V})}$ In this case we appeal to Lemma 4.2 to find that both $\mathcal{F}(V)$ and $\mathcal{F}(\widetilde{V})$ are completely reducible as $\Gamma/H$-graded modules with each direct summand isomorphic to a twist by some $f\in\widehat{\Gamma}/H^{\perp}$. In particular, $V$ appears as a direct summand of $\mathcal{F}(\widetilde{V})$, so we must have $V\cong\widetilde{V}^{f}$ for some $f\in\widehat{\Gamma}/H^{\perp}$. That is, $V$ and $\widetilde{V}$ are equivalent in $\mathsf{Irr}_{\Gamma/H}^{\textup{tw}}$ and thus $\mathcal{F}$ is injective.
For surjectivity, take an arbitrary finite dimensional $\Gamma$-graded irreducible module $W$. By Lemma 4.1 and Lemma 4.5, this module appears as a submodule of the loop module of some $\Gamma/H$-graded irreducible module $V$. By Theorem 4.6, the only $\Gamma$-graded submodules of $\mathrm{L}_{H}{(V)}$ are either $\mathrm{L}_{H}{(V)}$ itself or $V$ with a suitable $\Gamma$-grading. If $V$ can be given a $\Gamma$-grading then it appears as a $\Gamma$-graded submodule of $\mathrm{L}_{H}{(V)}$ (by 3.2) making $\mathrm{L}_{H}{(V)}$ reducible. Thus, there is only one option (up to $H$-parity shift) for what $W$ could have been. This only option is $\mathcal{F}(V)$ and so $\mathcal{F}$ is surjective. ∎
Recall that a similar bijection was noted in [61, Remark 7.2] via the use of Brauer invariants. We did not make use of Bauer invariants, but instead assumed that $H$ is a simple group. This assumption gave the bijection $\mathcal{F}$ the very simple form (4).
However, in general, it may be difficult to determine which case of $\mathcal{F}$ in (4) applies. There are multiple options we can test:
  • Try to find a $\Gamma$-grading for $V$ (see [56] for some results about $\Gamma$-gradability).
  • Try to find $\Gamma$-graded proper submodules of $\mathrm{L}_{H}{(V)}$ (all of which will provide a $\Gamma$-grading for $V$) or else prove that $\mathrm{L}_{H}{(V)}$ is $\Gamma$-graded irreducible (and hence $V$ will not be $\Gamma$-gradable).
  • Check whether the twists of $V$ by elements of $\widehat{\Gamma}/H^{\perp}$ are isomorphic. If not, then $V$ cannot be $\Gamma$-graded (by 4.3 (iii)).
Taking $\Gamma=\mathbb{Z}_{2}\times\mathbb{Z}_{2},\,H=\{00,11\}$, the above theorem (in combination with discolouration/recolouration) gives a way to construct the graded irreducible modules of the $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$-graded colour Lie superalgebras from the $\mathbb{Z}_{2}$-graded Lie superalgebras. More generally, we can use iterated applications of the theorem to produce all of the $\Gamma$-graded irreducible modules for any finite abelian group $\Gamma$. Theorem 4.8 also gives us a bijection between the graded and ungraded modules of the colour algebra.
5. Example: colour $\mathfrak{sl}_{2}$
As an example of applying Theorem 4.8, we will determine the finite-dimensional irreducible representations of $\mathfrak{sl}_{2}^{\mathrm{c}}$ over $\mathbb{C}$. The algebra $\mathfrak{sl}_{2}^{\mathrm{c}}$ is a $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$-graded Lie colour algebra with commutation factor $\varepsilon(\alpha_{1}\alpha_{2},\beta_{1}\beta_{2})=(-1)^{\alpha_{1}\beta_{2}-\alpha_{2}\beta_{1}}$ and sectors spanned by the following elements:
[…]
[$\commutator{a_{1}}{a_{2}}$ ] $=a_{3}$ [$\commutator{a_{2}}{a_{3}}$ ] $=-a_{1}$ [$\commutator{a_{3}}{a_{1}}$ ] $=-a_{2}.$
These commutation relations are satisfied by the following realisation
(5) $a_{1}$ $=\frac{i}{2}(e-f)$ $a_{2}$ $=-\frac{1}{2}(e+f)$ $a_{3}$ $=-\frac{i}{2}h.$
[…]
Both the ungraded [62] and graded [63] irreducible representations have appeared in the literature. In [62], the authors found all the ungraded irreducible representations for $\mathfrak{sl}_{2}^{\mathrm{c}}$ by embedding the universal enveloping algebra $U(\mathfrak{sl}_{2}^{\mathrm{c}})$ in $M_{2}(U(\mathfrak{sl}_{2}))$ (the space of $2\times 2$ matrices with entries in $U(\mathfrak{sl}_{2})$). In [63], the corresponding graded irreducible representations for the $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$-graded (non-colour) $\mathfrak{sl}_{2}$ were obtained using ideals generated by central elements of the universal enveloping algebra $U(\mathfrak{sl}_{2})$. The approach that we will take is presented in the more general context of the loop module construction.
Our strategy is as follows:
  1. (1)
    find all of the ungraded irreducible representations for $\mathfrak{sl}_{2}$;
  2. (2)
    use the techniques from the preceding sections to find the $\mathbb{Z}_{2}$-graded irreducible representations for $\mathfrak{sl}_{2}$;
  3. (3)
    similarly, find the $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$-graded irreducible representations for $\mathfrak{sl}_{2}$;
  4. (4)
    recolour to obtain the $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$-graded irreducible representations for $\mathfrak{sl}_{2}^{\mathrm{c}}$;
  5. (5)
    find the ungraded irreducible representations for $\mathfrak{sl}_{2}^{\mathrm{c}}$ as subrepresentations.
5.1. Irreducible representations for $\mathfrak{sl}_{2}$
Recall that $\mathfrak{sl}_{2}$ has a unique irreducible representation $V_{\lambda}$ of dimension $\lambda+1$ for every highest weight $\lambda\in\mathbb{Z}_{\geq 0}$.We will choose a basis $\{v_{0},v_{1},\ldots,v_{\lambda}\}$ for $V_{\lambda}$ such that
$hv_{j}$ $=(\lambda-2j)v_{j}$ $ev_{j}$ $=(\lambda-j+1)v_{j-1}$ $fv_{j}$ $=(j+1)v_{j+1}$
with the convention $v_{-1}=v_{\lambda+1}=0$. In particular, note that $v_{0}$ is the highest weight vector. Using (5), we find that
$a_{1}v_{j}$ $=\frac{i}{2}((\lambda-j+1)v_{j-1}-(j+1)v_{j+1})$
$a_{2}v_{j}$ $=-\frac{1}{2}((\lambda-j+1)v_{j-1}+(j+1)v_{j+1})$
$a_{3}v_{j}$ $=-\frac{i}{2}(\lambda-2j)v_{j}.$
5.2. $\mathbb{Z}_{2}$-graded irreducible representations
We wish to find the $\mathbb{Z}_{2}$-graded representations. So, we choose $\Gamma_{1}=H_{1}=\mathbb{Z}_{2}\times\mathbb{Z}_{2}/\{00,11\}\cong\mathbb{Z}_{2}$. Under this group and the original $\mathfrak{sl}_{2}^{\mathrm{c}}$ grading, the sectors for $\mathfrak{sl}_{2}$ are spanned by the following elements:
[…]
The ungraded module $V_{\lambda}$ can be given a $\Gamma_{1}$-grading: $V_{\lambda,0}^{\mathrm{E}}\oplus V_{\lambda,1}^{\mathrm{E}}$ where
$V_{\lambda,0}^{\mathrm{E}}$ $=\operatorname{span}\{v_{j}\mid j~{}\text{even}\}$ $V_{\lambda,1}^{\mathrm{E}}$ $=\operatorname{span}\{v_{j}\mid j~{}\text{odd}\}.$
We can easily check that $a_{1}V_{\lambda,\gamma}^{\mathrm{E}}\subseteq V_{\lambda,\gamma+1}^{\mathrm{E}},\,a_{2}V_{\lambda,\gamma}^{\mathrm{E}}\subseteq V_{\lambda,\gamma+1}^{\mathrm{E}}$ and $a_{3}V_{\lambda,\gamma}^{\mathrm{E}}\subseteq V_{\lambda,\gamma}^{\mathrm{E}}$ for $\gamma\in\mathbb{Z}_{2}$. Call this $\Gamma_{1}$-graded module $V_{\lambda}^{\mathrm{E}}$.
Since $V_{\lambda}$ is ungraded irreducible, $V_{\lambda}^{E}$ is $\Gamma_{1}$-graded irreducible. By shifting the $H_{1}$-parity of $V_{\lambda}^{\mathrm{E}}$, we get another $\Gamma_{1}$-graded irreducible module $V_{\lambda}^{\mathrm{O}}=V_{\lambda,0}^{\mathrm{O}}\oplus V_{\lambda,1}^{\mathrm{O}}$ where
$V_{\lambda,0}^{\mathrm{O}}$ $=\operatorname{span}\{v_{j}\mid j~{}\text{odd}\}$ $V_{\lambda,1}^{\mathrm{O}}$ $=\operatorname{span}\{v_{j}\mid j~{}\text{even}\}.$
By 4.7, we do not need to search for any other gradings. And by 3.2, we do not need to check the loop module at all. Thus, the finite-dimensional $\Gamma_{1}$-graded irreducible $\mathfrak{sl}_{2}$-modules are $V_{\lambda}^{\mathrm{E}}$ and $V_{\lambda}^{\mathrm{O}}$ for $\lambda\in\mathbb{Z}_{\geq 0}$.
5.3. $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$-graded irreducible representations
Now we wish to find the $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$-graded representations. So, we choose $\Gamma_{2}=\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ and $H_{2}=\{00,11\}$. We will now find the $\Gamma_{2}$-gradings, with the aid of the following lemma.
Lemma 5.1.
Let $V_{\lambda}=\bigoplus_{\gamma\in\Gamma_{2}}W_{\gamma}$ be a $\Gamma_{2}$-grading for $V_{\lambda}$. Then $v_{0}=\sum_{\gamma\in\Gamma_{2}}v_{0,\gamma}$ where $v_{0,\gamma}\in W_{\gamma}$ is some linear combination of $v_{0}$ and $v_{\lambda}$ (possibly equal to zero).
Proof.
Since $h=2ia_{3}$ is homogeneous of degree $11$, we find that $h^{2}W_{\gamma}\subseteq W_{\gamma}$. If we project $v_{0}$ onto each sector and write $v_{0}=\sum_{\gamma\in\Gamma_{2}}v_{0,\gamma}$, for $v_{0,\gamma}\in W_{\gamma}$ then we find that
$\sum_{\gamma\in\Gamma_{2}}h^{2}v_{0,\gamma}=h^{2}(v_{0})=\sum_{\gamma\in\Gamma_{2}}\lambda^{2}v_{0,\gamma}.$
In particular, $h^{2}v_{0,\gamma}=\lambda^{2}v_{0,\gamma}$ for every $\gamma\in\Gamma_{2}$. Using our knowledge of the eigenvectors of $h$, we find that $v_{0,\gamma}$ must be a linear combination of $v_{0}$ and $v_{\lambda}$. ∎
Remark 5.2.
A similar result holds for $v_{\lambda}=\sum_{\gamma\in\Gamma_{2}}v_{\lambda,\gamma}$.
Lemma 5.1 tells us that at least one sector of any $\Gamma_{2}$-grading must contain a nonzero linear combination of $v_{0}$ and $v_{\lambda}$. Our strategy is to then apply $a_{1}$ and $a_{2}$ to this linear combination to help deduce the structure of a $\Gamma_{2}$-grading or prove that one cannot exist.
5.3.1. The $\lambda$ even case
If $\lambda$ is even, then both $v_{0}$ and $v_{\lambda}$ are in $V_{\lambda,0}^{\mathrm{E}}$. Therefore, using Lemma 5.1, we find that the $00$-sector or $11$-sector must contain a nonzero linear combination of $v_{0}$ and $v_{\lambda}$. We choose (after some trial and error) $v_{0}+v_{\lambda}$ to be in the $00$-sector and $v_{0}-v_{\lambda}$ to be in the $11$-sector. With this choice, we can find a $\Gamma_{2}$-grading for $V_{\lambda}^{\mathrm{E}}$:
(6) $V_{\lambda,00}^{\mathrm{E}+}$ $=\operatorname{span}\{v_{j}+v_{\lambda-j}\mid j~{}\text{even}\},$ $V_{\lambda,10}^{\mathrm{E}+}$ $=\operatorname{span}\{v_{j}-v_{\lambda-j}\mid j~{}\text{odd}\},$
$V_{\lambda,01}^{\mathrm{E}+}$ $=\operatorname{span}\{v_{j}+v_{\lambda-j}\mid j~{}\text{odd}\},$ $V_{\lambda,11}^{\mathrm{E}+}$ $=\operatorname{span}\{v_{j}-v_{\lambda-j}\mid j~{}\text{even}\}.$
Denote this $\Gamma_{2}$-graded module by $V_{\lambda}^{\mathrm{E}+}$. We calculate
$a_{1}(v_{j}\pm v_{\lambda-j})$ $=\frac{i}{2}((\lambda-j+1)(v_{j-1}\mp v_{\lambda-(j-1)})-(j+1)(v_{j+1}\mp v_{\lambda-(j+1)})$
$a_{2}(v_{j}\pm v_{\lambda-j})$ $=-\frac{1}{2}((\lambda-j+1)(v_{j-1}\pm v_{\lambda-(j-1)})+(j+1)(v_{j+1}\pm v_{\lambda-(j+1)})$
$a_{3}(v_{j}\pm v_{\lambda-j})$ $=-\frac{i}{2}((\lambda-2j)(v_{j}\mp v_{\lambda-j}))$
from which we can easily verify that $a_{1}V_{\lambda,\gamma}^{\mathrm{E}+}\subseteq V_{\lambda,\gamma+10}^{\mathrm{E}+}$, $a_{2}V_{\lambda,\gamma}^{\mathrm{E}+}\subseteq V_{\lambda,\gamma+01}^{\mathrm{E}+}$ and $a_{3}V_{\lambda,\gamma}^{\mathrm{E}+}\subseteq V_{\lambda,\gamma+11}^{\mathrm{E}+}$ for each $\gamma\in\Gamma_{2}$. Since $V_{\lambda}^{\mathrm{E}}$ is $\Gamma_{2}$-graded irreducible, so is $V_{\lambda}^{\mathrm{E}+}$. By shifting the $H_{2}$-parity of $V_{\lambda}^{\mathrm{E}+}$, we get another $\Gamma_{2}$-graded irreducible module $V_{\lambda}^{\mathrm{E}-}$, which can be obtained from $V_{\lambda}^{\mathrm{E}+}$ by swapping the $00$- with the $11$-sector and the $10$- with the $01$-sector. Note that $V_{\lambda}^{\mathrm{E}+}$ and $V_{\lambda}^{\mathrm{E}-}$ are not isomorphic, because $v_{\lambda/2}=(1/2)(v_{\lambda/2}+v_{\lambda-\lambda/2})$ (the unique vector with weight 00) will be in different sectors.
Performing similar computations with $V_{\lambda}^{\mathrm{O}}$, we find two more $\Gamma_{2}$-graded irreducible modules $V_{\lambda}^{\mathrm{O}+}$ and $V_{\lambda}^{\mathrm{O}-}$ (these can be obtained from $V_{\lambda}^{\mathrm{E}+}$ by swapping the $00$- with the $01$-sector and the $10$- with the $11$-sector to get $V_{\lambda}^{\mathrm{O}+}$; and by swapping the $00$- with the $10$-sector and the $01$- with the $11$-sector to get $V_{\lambda}^{\mathrm{O}-}$). By 4.7 and Theorem 4.8 we do not need to search for any more $\Gamma_{2}$-graded irreducible modules in the case where $\lambda$ is even.
5.3.2. The $\lambda$ odd case
Lemma 5.3.
If $\lambda$ is odd, then $V_{\lambda}^{\mathrm{E}}$ cannot be given a $\Gamma_{2}$-grading.
Proof.
For contradiction, assume that we can give $V_{\lambda}^{\mathrm{E}}$ a $\Gamma_{2}$-grading, $V_{\lambda}^{\mathrm{E}}=\bigoplus_{\gamma\in\Gamma_{2}}V_{\lambda,\gamma}^{\mathrm{E}}$. From the structure of $V_{\lambda}^{\mathrm{E}}$, we know that $v_{0}\in V_{\lambda,0}^{\mathrm{E}}=V_{\lambda,00}^{\mathrm{E}}\oplus V_{\lambda,11}^{\mathrm{E}}$. Therefore, we can write $v_{0}=v_{0,00}+v_{0,11}$ for $v_{0,00}\in V_{\lambda,00}^{\mathrm{E}},\,v_{0,11}\in V_{\lambda,11}^{\mathrm{E}}$. By Lemma 5.1, both $v_{0,00}$ and $v_{0,11}$ are linear combinations of $v_{0}$ and $v_{\lambda}$. But since $v_{\lambda}\in V_{\lambda,1}^{\mathrm{E}}=V_{\lambda,10}^{\mathrm{E}}\oplus V_{\lambda,01}^{\mathrm{E}}$, we must have that both $v_{0,00}$ and $v_{0,11}$ are scalar multiples of $v_{0}$. And since they sum to $v_{0}\neq 0$, at least one of them must be nonzero. That is, $v_{0}\in V_{\lambda,00}^{\mathrm{E}}$ or $v_{0}\in V_{\lambda,11}^{\mathrm{E}}$. An analogous argument shows that $v_{\lambda}\in V_{\lambda,10}$ or $v_{\lambda}\in V_{\lambda,01}$.
If $v_{0}\in V_{\lambda,00}^{\mathrm{E}}$ and $v_{\lambda}\in V_{\lambda,01}^{\mathrm{E}}$ then applying $a_{1}$ we find
$v_{1}$ $=2ia_{1}v_{0}\in V_{\lambda,00}^{\mathrm{E}}\oplus V_{\lambda,10}^{\mathrm{E}}$
$v_{2}$ $=ia_{1}v_{1}+\frac{\lambda}{2}v_{0}\in V_{\lambda,00}^{\mathrm{E}}\oplus V_{\lambda,10}^{\mathrm{E}}$
[…]
$v_{j+1}$ $=\frac{2i}{j+1}a_{1}v_{j}+\frac{\lambda-j+1}{j+1}v_{j-1}\in V_{\lambda,00}^{\mathrm{E}}\oplus V_{\lambda,10}^{\mathrm{E}}.$
Proceeding inductively, we find that $V_{\lambda}^{\mathrm{E}}\subseteq V_{\lambda,00}^{\mathrm{E}}\oplus V_{\lambda,10}^{\mathrm{E}}$. But, by applying $a_{1}$ to $v_{\lambda}$ in a similar way, we find that $V_{\lambda}^{\mathrm{E}}\subseteq V_{\lambda,01}^{\mathrm{E}}\oplus V_{\lambda,11}^{\mathrm{E}}$, a contradiction! We similarly arrive at a contradiction if we assume $v_{0}\in V_{\lambda,11}^{\mathrm{E}}$ and $v_{\lambda}\in V_{\lambda,10}^{\mathrm{E}}$; $v_{0}\in V_{\lambda,00}^{\mathrm{E}}$ and $v_{\lambda}\in V_{\lambda,10}^{\mathrm{E}}$; or $v_{0}\in V_{\lambda,11}$ and $v_{\lambda}\in V_{\lambda,01}$. (Note that in some of these cases we need to use $a_{2}$ instead of $a_{1}$.) ∎
We have just shown that $V_{\lambda}^{\mathrm{E}}$ cannot be given a $\Gamma_{2}$-grading if $\lambda$ is odd. By Theorem 4.6, the loop module $\mathrm{L}_{H_{2}}{(V_{\lambda}^{\mathrm{E}})}$ must therefore be a $\Gamma_{2}$-graded irreducible representation. If we take a basis $\{v_{\alpha,j}\mid\alpha\in H_{2}\cong\mathbb{Z}_{2},\,j=0,\ldots,\lambda\}$ for $\mathrm{L}_{H_{2}}{(V_{\lambda}^{\mathrm{E}})}$, where
(7) $\mathrm{L}_{H_{2}}{(V^{\mathrm{E}}_{\lambda})}_{00}$ $=\operatorname{span}\{v_{0,j}\mid j~{}\text{even}\},$ $\mathrm{L}_{H_{2}}{(V^{\mathrm{E}}_{\lambda})}_{10}$ $=\operatorname{span}\{v_{1,j}\mid j~{}\text{odd}\},$
$\mathrm{L}_{H_{2}}{(V^{\mathrm{E}}_{\lambda})}_{01}$ $=\operatorname{span}\{v_{0,j}\mid j~{}\text{odd}\},$ $\mathrm{L}_{H_{2}}{(V^{\mathrm{E}}_{\lambda})}_{11}$ $=\operatorname{span}\{v_{1,j}\mid j~{}\text{even}\},$
then we can compute the action as
$a_{1}v_{\alpha,j}$ $=\frac{i}{2}((\lambda-j+1)v_{\alpha+1,j-1}-(j+1)v_{\alpha+1,j+1})$
$a_{2}v_{\alpha,j}$ $=-\frac{1}{2}((\lambda-j+1)v_{\alpha,j-1}+(j+1)v_{\alpha,j+1})$
$a_{3}v_{\alpha,j}$ $=-\frac{i}{2}(\lambda-2j)v_{\alpha+1,j}.$
By similar reasoning, we find that $\mathrm{L}_{H_{2}}{(V_{\lambda}^{\mathrm{O}})}$ is a $\Gamma_{2}$-graded irreducible representation. We can obtain $\mathrm{L}_{H_{2}}{(V_{\lambda}^{\mathrm{O}})}$ from $\mathrm{L}_{H_{2}}{(V_{\lambda}^{\mathrm{E}})}$ by swapping the $00$- with the $01$-sector and the $10$- with the $11$-sector. Note that if we shift the $H_{2}$-parity of either of these modules, we get the same module. By Theorem 4.8, the only $\Gamma_{2}$-graded irreducible modules for the case where $\lambda$ is odd are $\mathrm{L}_{H_{2}}{(V_{\lambda}^{\mathrm{E}})}$ and $\mathrm{L}_{H_{2}}{(V_{\lambda}^{\mathrm{O}})}$.
5.4. Recolouring
To recolour $\mathfrak{sl}_{2}$ to $\mathfrak{sl}_{2}^{\mathrm{c}}$ we use the multiplier $\sigma(\alpha_{1}\alpha_{2},\beta_{1}\beta_{2})=(-1)^{\alpha_{2}\beta_{1}}$. To obtain the $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$-graded irreducible modules for $\mathfrak{sl}_{2}^{\mathrm{c}}$, we use this same multiplier for the $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$-graded irreducible modules for $\mathfrak{sl}_{2}$.
Theorem 5.4.
The only finite-dimensional $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$-graded irreducible modules for $\mathfrak{sl}_{2}^{\mathrm{c}}$ are:
  1. (i)
    For each even number $\lambda$, the four modules which are equivalent to $(V_{\lambda}^{\mathrm{E}+})^{\sigma}$ up to $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$-parity shift. The sectors of $(V_{\lambda}^{\mathrm{E}+})^{\sigma}$ are given as in (6) and the action is given by
    $a_{1}(v_{j}\pm v_{\lambda-j})$ $=\frac{i}{2}((\lambda-j+1)(v_{j-1}\mp v_{\lambda-(j-1)})-(j+1)(v_{j+1}\mp v_{\lambda-(j+1)})$
    $a_{2}(v_{j}\pm v_{\lambda-j})$ $=\mp\frac{1}{2}((\lambda-j+1)(v_{j-1}\pm v_{\lambda-(j-1)})+(j+1)(v_{j+1}\pm v_{\lambda-(j+1)})$
    $a_{3}(v_{j}\pm v_{\lambda-j})$ $=\mp\frac{i}{2}((\lambda-2j)(v_{j}\mp v_{\lambda-j})).$
    The dimension of these modules is $\lambda+1$.
  2. (ii)
    For each odd number $\lambda$, the two modules which are equivalent to $(\mathrm{L}_{H_{2}}{(V_{\lambda}^{\mathrm{E}})})^{\sigma}$ up to $\{00,01\}$-parity shift. The sectors of $(\mathrm{L}_{H_{2}}{(V_{\lambda}^{\mathrm{E}})})^{\sigma}$ are given as in (7) and the action is given by
    $a_{1}v_{\alpha,j}$ $=\frac{i}{2}((\lambda-j+1)v_{\alpha+1,j-1}-(j+1)v_{\alpha+1,j+1})$
    $a_{2}v_{\alpha,j}$ $=-\frac{(-1)^{\alpha}}{2}((\lambda-j+1)v_{\alpha,j-1}+(j+1)v_{\alpha,j+1})$
    $a_{3}v_{\alpha,j}$ $=-\frac{i(-1)^{\alpha}}{2}(\lambda-2j)v_{\alpha+1,j}.$
    The dimension of these modules is $2(\lambda+1)$.
Proof.
Use the results from the previous sections and the facts that $\sigma(10,\beta)=1$ for all $\beta\in\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ and
$\sigma(01,\beta)=\sigma(11,\beta)=\begin{cases}-1&\text{if}~{}\beta=10,11\\ 1&\text{if}~{}\beta=00,01.\end{cases}\qed$
Note that the irreducible modules in the above theorem are recoloured versions of the torsion-free modules previously appearing in [63]. The above theorem proves, in addition, that these modules are the only $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$-graded irreducible $\mathfrak{sl}_{2}^{\mathrm{c}}$ modules.
Note that, for each $\lambda\in\mathbb{Z}_{\geq 0}$, there is a unique equivalence class of $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$-graded irreducible $\mathfrak{sl}_{2}^{\mathrm{c}}$-modules. This corresponds exactly to the unique irreducible $\mathfrak{sl}_{2}$-modules, demonstrating the bijection in Theorem 4.8.
5.5. Ungraded irreducible representations for $\mathfrak{sl}_{2}^{\mathrm{c}}$
Lemma 5.5.
The ungraded irreducible representations for $\mathfrak{sl}_{2}^{\mathrm{c}}$ all appear as ungraded subrepresentations of the $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$-graded irreducible representations in Theorem 5.4.
Proof.
If we knew the ungraded irreducible representations for $\mathfrak{sl}_{2}^{\mathrm{c}}$, then we could use the same process in the preceding sections to derive the same $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$-graded irreducible modules. By applying Theorem 4.8 twice, these $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$-graded irreducible modules are either irreducible as ungraded modules, a loop module or a loop module of a loop module. In any case, the original ungraded irreducible modules must appear as ungraded submodules of the $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$-graded modules. ∎
Note that $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$-graded modules which are equivalent up to $H$-parity shift (for some $H\leq\mathbb{Z}_{2}$) are isomorphic as ungraded modules.
Lemma 5.6.
If $\lambda$ is even, then $(V_{\lambda}^{E+})^{\sigma}$ is irreducible as an ungraded module.
Proof.
Choose a new basis $\{u_{j}\mid j=0,\ldots,\lambda\}$ where $u_{j}=(v_{j}+v_{\lambda-j})+i(-1)^{j}(v_{j}-v_{\lambda-j})$. Note that this is not linearly dependent since $u_{\lambda-j}=(v_{j}+v_{\lambda-j})-i(-1)^{j}(v_{j}-v_{\lambda-j})$. The action of $\mathfrak{sl}_{2}^{\mathrm{c}}$ is then
(8) $a_{1}u_{j}$ $=\frac{(-1)^{j+1}}{2}((\lambda-j+1)u_{j-1}-(j+1)u_{j+1})$
$a_{2}u_{j}$ $=-\frac{1}{2}((\lambda-j+1)u_{j-1}+(j+1)u_{j+1})$
$a_{3}u_{j}$ $=\frac{(-1)^{j+1}}{2}(\lambda-2j)u_{j}$
with the convention that $u_{-1}=u_{\lambda+1}=0$. In particular,
$(a_{2}+a_{1})u_{j}$ =$=\begin{cases}-(\lambda-j+1)u_{j-1}&\text{if $j$ even}\\ -(j+1)u_{j+1}&\text{if $j$ odd}.\end{cases}$
$(a_{2}-a_{1})u_{j}$ =$=\begin{cases}-(j+1)u_{j+1}&\text{if $j$ even}\\ -(\lambda-j+1)u_{j-1}&\text{if $j$ odd}.\\ \end{cases}$
So, if we take an arbitrary element $\sum_{j}c_{j}u_{j}$ of $(V_{\lambda}^{E+})^{\sigma}$, then we can alternate applying $(a_{2}+a_{1})$ and $(a_{2}-a_{1})$ (i.e. we apply $((a_{2}-a_{1})(a_{2}+a_{1}))^{k}$ or $(a_{2}-a_{1})((a_{2}+a_{1})(a_{2}-a_{1}))^{k}$ for some non-negative integer $k$) to raise/lower the basis elements of $\sum_{j}c_{j}u_{j}$. Eventually, this procedure will yield a scalar multiple of either $u_{0}$ or $u_{\lambda}$, from which we can generate the entire module. Since the entire module can be generated from a single vector, $(V_{\lambda}^{\mathrm{E}+})^{\sigma}$ is irreducible. ∎
Now we consider the case when $\lambda$ is odd. Then $(\mathrm{L}_{H_{2}}{(V_{\lambda}^{\mathrm{E}})})^{\sigma}=U_{\lambda}^{++}\oplus U_{\lambda}^{+-}\oplus U_{\lambda}^{-+}\oplus U_{\lambda}^{--}$ where
$U_{\lambda}^{\zeta\xi}$ $=\operatorname{span}\{u^{\zeta\xi}_{j}\mid j=0,1,\ldots,(\lambda-1)/2\}$
$u^{\zeta\xi}_{j}$ $=v_{0,j}+\zeta i(-1)^{j}v_{1,j}+\xi v_{0,\lambda-j}-\zeta\xi i(-1)^{j}v_{1,\lambda-j}$
for $\zeta,\xi\in\{-1,+1\}$. The action of $\mathfrak{sl}_{2}^{\mathrm{c}}$ is given by
(9) $a_{1}u^{\zeta\xi}_{j}$ $=-\frac{\zeta}{2}(-1)^{j}((\lambda-j+1)u^{\zeta\xi}_{j-1}-(j+1)u^{\zeta\xi}_{j+1})$ $\left(j<\frac{\lambda-1}{2}\right)$
$a_{1}u^{\zeta\xi}_{(\lambda-1)/2}$ $=-\frac{\zeta}{2}(-1)^{(\lambda-1)/2}\left(\frac{\lambda+3}{2}u^{\zeta\xi}_{(\lambda-3)/2}-\xi\frac{\lambda+1}{2}u^{\zeta\xi}_{(\lambda-1)/2}\right)$
$a_{2}u^{\zeta\xi}_{j}$ $=-\frac{1}{2}((\lambda-j+1)u^{\zeta\xi}_{j-1}+(j+1)u^{\zeta\xi}_{j+1})$ $\left(j<\frac{\lambda-1}{2}\right)$
$a_{2}u^{\zeta\xi}_{(\lambda-1)/2}$ $=-\frac{1}{2}\left(\frac{\lambda+3}{2}u^{\zeta\xi}_{(\lambda-3)/2}+\xi\frac{\lambda+1}{2}u^{\zeta\xi}_{(\lambda-1)/2}\right)$
$a_{3}u^{\zeta\xi}_{j}$ $=-\frac{\zeta}{2}(-1)^{j}(\lambda-2j)u^{\zeta\xi}_{j}$
with the convention $u_{-1}=1$. (Note that $u^{\zeta\xi}_{j}$ was chosen so that it was an eigenvector of $a_{3}$.) Clearly $U_{\lambda}^{\zeta\xi}$ gives rise to four non-isomorphic modules, since the action of $a_{1}$, $a_{2}$ and $a_{3}$ on $u^{\zeta\xi}_{(\lambda-1)/2}$ (the unique eigenvector for $a_{3}$ which has eigenvalue with absolute value $1/2$) is different for different values of $\zeta,\xi$.
Lemma 5.7.
$U_{\lambda}^{\zeta\xi}$ is irreducible as an ungraded module.
Proof.
Observe that, for $j<(\lambda-1)/2$,
$(a_{2}+a_{1})u^{\zeta\xi}_{j}$ =$=\begin{cases}-(\lambda-j+1)u^{\zeta\xi}_{j-1}&\text{if}~{}\zeta(-1)^{j}=1\\ -(j+1)u^{\zeta\xi}_{j+1}&\text{if}~{}\zeta(-1)^{j}=-1\end{cases}$
$(a_{2}+a_{1})u^{\zeta\xi}_{(\lambda-1)/2}$ =$=\begin{cases}-\frac{\lambda+3}{2}u^{\zeta\xi}_{(\lambda-3)/2}&\text{if}~{}\zeta(-1)^{(\lambda-1)/2}=1\\ -\xi\frac{\lambda+1}{2}u^{\zeta\xi}_{(\lambda-1)/2}&\text{if}~{}\zeta(-1)^{(\lambda-1)/2}=-1\end{cases}$
$(a_{2}-a_{1})u^{\zeta\xi}_{j}$ =$=\begin{cases}-(j+1)u^{\zeta\xi}_{j+1}&\text{if}~{}\zeta(-1)^{j}=1\\ -(\lambda-j+1)u^{\zeta\xi}_{j-1}&\text{if}~{}\zeta(-1)^{j}=-1\end{cases}$
$(a_{2}-a_{1})u^{\zeta\xi}_{(\lambda-1)/2}$ =$=\begin{cases}-\xi\frac{\lambda+1}{2}u^{\zeta\xi}_{(\lambda-1)/2}&\text{if}~{}\zeta(-1)^{(\lambda-1)/2}=1\\ -\frac{\lambda+3}{2}u^{\zeta\xi}_{(\lambda-3)/2}&\text{if}~{}\zeta(-1)^{(\lambda-1)/2}=-1\end{cases}$
Take an arbitrary element $\sum_{j}c_{j}u_{j}$ of $U_{\lambda}^{\zeta\xi}$. Similar to the proof of Lemma 5.6 we alternate applying $a_{2}+a_{1}$ and $a_{2}-a_{1}$. In this case, we are able to obtain a scalar multiple of $u_{0}$ from $\sum_{j}c_{j}u_{j}$, from which we can generate the entire module. ∎
In summary, we have proven the following theorem:
Theorem 5.8.
The only finite-dimensional ungraded irreducible modules for $\mathfrak{sl}_{2}^{\mathrm{c}}$ are
  1. (i)
    For each even number $\lambda$, the module $(V_{\lambda}^{\mathrm{E}+})^{\sigma}$ with action given in (8). The dimension of these modules is $\lambda+1$.
  2. (ii)
    For each odd number $\lambda$ and each $\zeta,\xi\in\{-1,+1\}$, the module $U_{\lambda}^{\zeta\xi}$ with action given in (9). The dimension of these modules is $(\lambda+1)/2$.
This theorem agrees with the previous result of [62, Theorem 5.2], but is presented in the context of the theory in this paper.
Remark 5.9.
It is clear that the modules $U_{\lambda}^{+\xi}$ and $U_{\lambda}^{-\xi}$ are twists of each other, with twisting character $f\in\widehat{\Gamma}/H^{\perp}$ given by $f(00)=f(01)=1$ and $f(10)=f(11)=-1$ (here we can choose $\Gamma=\mathbb{Z}_{2}\times\mathbb{Z}_{2}/\{00,01\}$ and $H=\Gamma$). If we instead choose $f(00)=f(11)=1$ and $f(01)=f(10)=-1$ (with $\Gamma=\mathbb{Z}_{2}\times\mathbb{Z}_{2}/\{00,11\},\,H=\Gamma$), then we find that the twist of $U_{\lambda}^{\zeta+}$ by $f$ is isomorphic to $U_{\lambda}^{\zeta-}$.
The above remark shows that $U^{\zeta\xi}$ are all equivalent up to twist by $\widehat{\Gamma}/H^{\perp}$ for different values of $\zeta,\xi$. Consequently, there is a unique equivalence class in Theorem 5.8 for each $\lambda\in\mathbb{Z}_{\geq 0}$. This corresponds exactly to the equivalence classes for $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$-graded $\mathfrak{sl}_{2}^{\mathrm{c}}$-modules in Theorem 5.4, again demonstrating the bijection in Theorem 4.8.
Remark 5.10.
Although the ungraded $\mathfrak{sl}_{2}^{\mathrm{c}}$ representations were derived from the ungraded $\mathfrak{sl}_{2}$ representations, they have a remarkably different structure. This shows that, despite the procedure outlined in this paper, colour algebras still have an interesting representation theory.
6. Conclusions
The loop module is a natural structure to consider in the study of Lie colour algebras, including their applications to quantum mechanical systems. In this direction, we showed that the Hilbert space of the $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$-graded quantum mechanical system of [23] can be obtained via a loop module. The loop module could potentially be a useful tool for constructing quantum mechanical systems in future work.
The main result of this paper (Theorem 4.8) is the existence of a bijection between finite-dimensional $\Gamma$- and $\Gamma/H$-graded irreducible representations (up to equivalence relations stronger than isomorphism). This bijection makes use of iterated loop modules following the simple groups in the Jordan–Hölder decomposition. A similar bijection had previously been noted in [61, Remark 7.2], but the iterated loop module by simple groups simplifies the construction. In combination with discolouration/recolouration, this bijection provides a general procedure for generating all of the $\Gamma$-graded Lie colour algebra representations from the well-studied $\mathbb{Z}_{2}$-graded Lie superalgebra representations.
As an example of applying the bijection, we constructed all of the irreducible representations for the Lie colour algebra $\mathfrak{sl}_{2}^{\mathrm{c}}$. These irreducible representations agree with those previously derived in [62, 63] using a different construction. Despite the bijection of Theorem 4.8, these $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$-graded Lie colour algebra representations have a remarkably different structure from those of the ungraded discoloured Lie algebra.
The bijection (Theorem 4.8) also has applications to the study of Lie colour algebra quantum mechanical systems. For example, we could potentially apply this procedure to construct the irreducible representations of $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$-supersymmetry algebras with $\mathcal{N}\geq 2$, whose classification is desired for the quantum mechanical systems studied in [33]. Future research should work to obtain a general classification of irreducible Lie colour algebra representations, possibly by applying the bijection of this paper to existing classifications.
Acknowledgements
The author would like to thank Kenneth Price for making him aware of [59], and an anonymous reviewer for directing him to the loop module construction of [58]. The author is especially grateful to Phil Isaac for his invaluable guidance, and for taking the time to suggest improvements to the manuscript.